Stock R has a beta of 2.2, Stock S has a beta of 0.4, the required return on an average stock is 11%, and the risk-free rate of return is 5%. By how much does the required return on the riskier stock exceed the required return on the less risky stock? Round your answer to two decimal places.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

Required return for:

R=5+(11-5)*2.2=18.2%

S=5+(11-5)*0.4=7.4%

Hence difference=18.2-7.4

=10.8%